# Occlusion-Aware Consistent Model Predictive Control for Robot Navigation in Occluded Obstacle-Dense Environments

### v4

@@ Abstract
Ensuring safety and motion consistency for robot navigation in occluded, obstacle-dense environments is a critical challenge. In this context, this study presents an occlusion-aware Consistent Model Predictive Control (CMPC) strategy. To account for the occluded obstacles, it incorporates adjustable risk regions that represent their potential future locations. Subsequently, dynamic risk boundary constraints are developed online to ensure safety. The CMPC then constructs multiple locally optimal trajectory branches (each tailored to different risk regions) to strike a balance between safety and performance. A shared consensus segment is generated to ensure smooth transitions between branches without significant velocity fluctuations, further preserving motion consistency. To facilitate high computational efficiency and ensure coordination across local trajectories, we use the alternating direction method of multipliers (ADMM) to decompose the CMPC into manageable sub-problems for parallel solving. The proposed strategy is validated through simulations and real-world experiments on an Ackermann-steering robot platform. The results demonstrate the effectiveness of the proposed CMPC strategy through comparisons with baseline approaches in occluded, obstacle-dense environments.
@@ I Introduction
Ensuring safe navigation and consistent motion for mobile robots in occluded, obstacle-dense environments is a critical challenge [1]. One of the key underlying factors to this concern is the partial observability of such environments due to occlusions [2, 3]. In this context, robots relying on onboard perception, which typically operate on line-of-sight principles, are unable to accurately detect occluded obstacles [4, 5]. This limitation increases the risk of collisions, particularly in obstacle-dense environments. Additionally, the sudden emergence of occluded obstacles can lead to abrupt velocity changes for the robot, compromising motion consistency and stability.
@@
Figure 1: A robot navigates an occluded, obstacle-dense environment with occluded dynamic obstacles. The planner generates three trajectories (orange, red, and blue) with different considerations of risk regions. All trajectories share an initial consensus segment (green) to enable smooth transitions between trajectories and ensure motion consistency.
@@
To ensure safe navigation in obstacle-dense environments, model predictive control (MPC) facilitates obstacle avoidance through receding-horizon optimization, which systematically accounts for system dynamics and environmental constraints to achieve proactive collision avoidance [6, 7]. To construct obstacle avoidance constraints, accurate prediction of obstacle states is essential for this strategy. Hence, robots must account for potential states of occluded obstacles to guarantee safety in occluded scenarios [1]. However, precise state estimation of fully occluded obstacles remains infeasible [8]. An alternative way involves predicting the probabilistic risk regions of these obstacles, which reduces planning complexity and enhances safety margins [4]. For instance, OA-MPC [1] adopts worst-case assumptions to ensure safety. This approach theoretically covers potential states of occluded obstacles. However, the generated trajectory tends to be overly conservative [9]. Moreover, a large deviation from the reference path compromises motion consistency [10].
@@
To enhance motion consistency, recent studies have proposed scenario-adaptive planning strategies [11, 12]. In these approaches, multiple homotopic local optimal trajectories are generated for autonomous vehicles. Although consistent parameters are used within the decision-making module to select the optimal trajectory and maintain motion consistency, each trajectory is optimized separately. Essentially, the lack of coordination can lead to velocity fluctuations, particularly in occluded, obstacle-dense settings. Alternatively, branch MPC (BMPC) [13, 14] introduces a shared common consensus segment across branches, ensuring smooth transitions between trajectory branches and maintaining motion consistency. Similarly, partially observable Markov decision processes (POMDP) [15] ensures environmental consistency through belief state updates [16]. However, these approaches encounter scalability challenges, as branch proliferation significantly increases the number of optimization variables, leading to intractable computation. To address this challenge, recent works have employed the alternating direction method of multipliers (ADMM), which decomposes the original problem into low-dimensional sub-problems to enable real-time performance without compromising safety [17, 18, 19]. Nonetheless, despite the strong computational performance of ADMM, a more definite strategy is still lacking on how to extend its usage to address the challenges posed by occluded, obstacle-dense environments, particularly in the real-world setting.
@@
In this study, we introduce a novel occlusion-aware Consistent Model Predictive Control (CMPC) strategy for mobile robots navigating in occluded, obstacle-dense environments. The CMPC optimizes several local optimal trajectories in parallel to account for potential occlusion risks, as shown in Fig. 1. Subsequently, an ADMM-based optimization strategy is used to decompose the optimization problem into parallel sub-problems to facilitate computational efficiency. The main contributions of this paper are summarized as follows:
@@ II Related Work
In this section, we review three major groups of approaches to occlusion-aware planning: reachability analysis, POMDP, and contingency planning.
@@ II-A Reachability Analysis
Reachability analysis is an efficient approach in occlusion-aware planning, as demonstrated in extensive studies [20, 4, 21, 22]. It creates a reachable set of potential agents in the occluded regions to construct constraints for planning. In [23], the impact of environmental visibility on mobile robot navigation is considered to ensure safety. An MPC strategy is proposed to apply risk fields from reachable sets for automated vehicles [24]. However, reachability analysis typically focuses on worst-case scenarios, leading to overly conservative behaviors. Additionally, predicting obstacle states based on rules of on-road scenarios [20, 4] is not applicable to occluded, obstacle-dense environments encountered by mobile robots.
@@ II-B POMDP
POMDP has been widely adopted in occlusion-aware planning for its ability to handle partially observable scenarios [15, 25]. It leverages a probability distribution in the prediction of the agents’ future states for effective planning. For instance, works [26, 27] leverage the POMDP framework to address navigation and decision-making problems in dynamic environments with partial occlusions. [16] further incorporates a belief state updating module to predict the scenarios more precisely for effective planning. While these approaches can tackle partially observed obstacles in structured urban road environments, their high computational complexity poses challenges for real-time planning. Additionally, obtaining reliable probability distributions is difficult in unstructured environments with dense occlusion, potentially leading to planning failures.
@@ II-C Contingency Planning
Contingency planning enhances safety and motion consistency by providing multiple contingency paths to address uncertainties in the future states of dynamic obstacles [28, 29, 13, 30]. [31] presents a learning-based method for predicting the states of partially observed agents and motion planning on road. However, existing approaches typically require substantial computational resources to determine policies based on assumptions about future states. Additionally, contingency planning typically makes rule-based assumptions about the future states in road scenarios, or one-to-one interaction between agents [10]. These future state assumptions may not be reliable in general scenarios.
In this study, we follow a simplified concept of reachability analysis to model the risk regions where occluded obstacles may appear in the future. We predict multiple configurations of risk regions while avoiding explicit hypotheses about obstacle states. To address computational complexity, we propose an occlusion-aware CMPC strategy combined with an ADMM-based optimization process. This strategy ensures safety, efficiency, and adaptability in occluded, obstacle-dense environments.
@@ III Problem Statement
where $\Delta t$ denotes the discrete time step; and $k$ denotes the time step. $f(s(k),u(k),\Delta t)$ is defined as follows:
@@
When an obstacle’s center is not visible in the robot’s FoV, it is treated as an occluded obstacle with unknown states.
@@
The objective of this work is to develop a motion planning strategy that generates a set of trajectories, enabling safe and consistent navigation in occluded, obstacle-dense environments. The proposed strategy must satisfy the following criteria:
Computational efficiency: Ensure real-time parallel trajectory generation in occluded, obstacle-dense environments.
Safety guarantee: Guarantee collision-free navigation in occluded, obstacle-dense environments.
@@ IV-A Risk Regions Modeling
To address the uncertainty of the states of occluded obstacles, we dynamically model risk regions to represent potential locations of occluded obstacles. The risk regions are continuously updated based on the latest states of the robot and obstacles.
@@ IV-A1 Definition of Occluded Regions
Occluded regions are defined as areas blocked by obstacles within the robot’s FoV. In the robot’s body frame, with its center as the origin, occluded regions are defined as the set of points that satisfy the following conditions:
@@
Based on this definition, the occluded region is described as the area bounded by two tangent lines of the visible obstacle. Let $m_{1}$ and $m_{2}$ denote the slopes of the two tangent lines that form the boundaries of the occluded region, $x_{\text{rel}}$ and $y_{\text{rel}}$ denote the position of the obstacle in the robot’s body frame, and $r_{\text{obs}}$ denotes the radius of the obstacle. The expressions for the two tangent lines are defined as follows:
@@
The occluded region $\mathcal{C}$ is represented by the set of points $\mathbf{x}=\begin{bmatrix}x\\ y\end{bmatrix}$ in the 2D plane that satisfy the following condition:
$\mathcal{C}=\{\mathbf{x}\in\mathbb{R}^{2}|\mathbf{A}\mathbf{x}<0\},$ (6)
where $\mathbf{A}=\begin{bmatrix}-m_{1}&1\\ m_{2}&-1\\ -1&0\end{bmatrix}$ denotes the coefficient matrix that represents the boundaries of the occluded region.
Figure 2: Modeling of occluded regions and risk regions. The occluded region is bounded by two tangent lines. Risk regions are located along two tangent lines, each with center $(c_{x},c_{y})$ and radius $r_{\text{risk}}$.
@@ IV-A2 Definition of Risk Regions
Based on the defined occluded regions, we further delineate several risk regions that affect the robot’s motion. Typically, obstacles at the edge of the occluded regions pose the greatest safety threat to the robot. Therefore, these risk regions are defined along the two tangent lines that form the occluded region described in Section IV-A1. Specifically, to provide an effective geometric approximation of the polygonal occlusion cones, we model two circular risk regions on each tangent line, as shown in Fig. 2. For the $i$-th circular risk region, the center $(c_{x}^{(i)},c_{y}^{(i)})$ is defined as follows:
$c_{x}^{(i)}$ $=p_{x}+(id_{\text{risk}}+\sqrt{d_{\text{obs}}^{2}-r_{\text{obs}}^{2}})\cos{(\arctan{m_{j}})},$ (7a)
$c_{y}^{(i)}$ $=p_{y}+(id_{\text{risk}}+\sqrt{d_{\text{obs}}^{2}-r_{\text{obs}}^{2}})\sin{(\arctan{m_{j}})},$ (7b)
$j$ $\in\mathcal{I}_{1}^{2},\ i\in\mathcal{I}_{0}^{N_{\text{risk}}},$
where $m_{j}$ denotes the slope of the $j$-th tangent line of the occluded region; $d_{\text{risk}}$ denotes the predefined distance between consecutive risk regions; $d_{\text{obs}}$ denotes the distance between the robot and the obstacle; $N_{\text{risk}}$ denotes the total number of risk regions in each configuration; and $\mathcal{I}_{a}^{b}$ denotes a sequence of integers from $a$ to $b$.
@@
$r_{\text{risk}}^{(i)}=\frac{\sqrt{(c_{x}^{(i)})^{2}+(c_{y}^{(i)})^{2}}}{v}\cdot v_{\text{obs,max}}+r_{\text{obs}},$ (8)
where $v$ denotes the velocity of the robot; and $v_{\text{obs,max}}$ denotes the assumed maximum velocity of the occluded obstacle.
In this study, the conservatism in trajectory planning can be adjusted by making different assumptions about the maximum velocity $v_{\text{obs, max}}$ to generate a sequence of $N_{z}$ risk region configurations.
@@
By utilizing the maximum possible obstacle velocity $v_{\text{obs,max}}$ in (8), our circular risk region model provides a conservative over-approximation of all potential positions that an occluded obstacle could occupy. This approach guarantees that the risk region encompasses any location an occluded obstacle may reach during the prediction horizon. For non-circular obstacles, we similarly use their minimum bounding circles to define the risk regions, ensuring the generality of our formulation.
@@ IV-B Occlusion-Aware CMPC
To complete the navigation task and enhance motion consistency, we define a cost function $\mathcal{J}$ to address these requirements. The cost function is defined over a prediction horizon of NN as follows:
@@
where $C_{\text{acc}}(\cdot)$ regulates the robot’s acceleration; $C_{\text{vel}}(\cdot)$ represents the velocity tracking cost, measured by the velocity error between the robot and reference velocity; and $C_{\text{guide}}(\cdot)$ represents the deviation from the guidance points. $w_{\text{acc}},\ w_{\text{vel}},\ w_{\text{guide}}\in\mathbb{R}^{+}$ are the corresponding weighting matrices for the cost components.
@@
A set of guidance points is generated using a Visual-PRM approach [33] for each horizon. It produces optimal guidance paths for the receding horizon based on a global task path, considering all visible obstacles. Specifically, guidance points provide only the endpoints of each horizon, so the corresponding cost $C_{\text{guide}}(\cdot)$ is only calculated at the time step k=N1k=N-1.
@@
$s(k)=\tilde{s}(k),\ \forall k\in\mathcal{I}_{0}^{N_{c}-1},$ (10e)
@@
$H_{\text{kin}}(s)=s(k+1)-f(s(k),u(k),\Delta t),\ \forall k\in\mathcal{I}_{0}^{N-1},$ (11a)
@@
$\forall\{(x_{\text{obs}},y_{\text{obs}}),r_{\text{obs}}\}\in\{O^{(j)}\}_{j=0}^{N_{\text{obs}}},\ \forall s\in\mathcal{S},$ (11b)
@@
$\forall\{(c_{x}^{(i)},c_{y}^{(i)}),r_{\text{risk}}^{(i)}\}\in\{R^{(i)}_{z}\}_{i=0}^{N_{\text{risk}}},\ \forall s\in\mathcal{S}.$ (11c)
Here, $H_{\text{kin}}(\cdot)$ in (11a) enforces kinematic constraints derived from the robot’s kinematic model defined in (3); $G_{\text{obs}}(\cdot)$ in (11b) ensures a minimum safety distance $r_{\text{obs}}$ between the robot and obstacle centers to avoid collision; $O^{(j)}$ represents the state of the $j$-th obstacle, with the position $(x_{\text{obs}},y_{\text{obs}})$ and the radius of the obstacle $r_{\text{obs}}$; $N_{\text{obs}}$ denotes the number of obstacles considered; $G_{\text{risk}}(\cdot)$ in (11c) maintains a risk boundary $r_{\text{risk}}^{(i)}$ around risk regions; $R^{(j)}_{(z)}$ represents the states of the $i$-th risk region in the $z$-th configuration, with center and radius derived from (7) and (8); $N_{z}$ denotes the total number of risk region configurations; (10e) represents the consensus constraint, with consensus variable $\tilde{s}$; and $N_{c}$ denotes the length of the consensus segment.

### v5

@@ Abstract
Ensuring safety and motion consistency for robot navigation in occluded, obstacle-dense environments is a critical challenge. In this context, this study presents an occlusion-aware Consistent Model Predictive Control (CMPC) strategy. To account for the occluded obstacles, it incorporates adjustable risk regions that represent their potential future locations. Subsequently, dynamic risk boundary constraints are developed online to enhance safety. Based on these constraints, the CMPC constructs multiple locally optimal trajectory branches (each tailored to different risk regions) to strike a balance between safety and performance. A shared consensus segment is generated to ensure smooth transitions between branches without significant velocity fluctuations, preserving motion consistency. To facilitate high computational efficiency and ensure coordination across local trajectories, we use the alternating direction method of multipliers (ADMM) to decompose the CMPC into manageable sub-problems for parallel solving. The proposed strategy is validated through simulations and real-world experiments on an Ackermann-steering robot platform. The results demonstrate the effectiveness of the proposed CMPC strategy through comparisons with baseline approaches in occluded, obstacle-dense environments.
@@ I Introduction
Ensuring safe navigation and consistent motion for mobile robots in occluded, obstacle-dense environments is a critical challenge [8]. One of the key underlying factors to this concern is the partial observability of such environments due to occlusions [1, 19]. In this context, robots relying on onboard perception, which typically operate on line-of-sight principles, are unable to detect obstacles hidden by occlusions [24, 2]. This limitation increases the risk of collisions, particularly in obstacle-dense environments. Additionally, the sudden emergence of occluded obstacles can lead to abrupt velocity changes for the robot, compromising motion consistency and stability.
@@
Figure 1: Robot navigation in an occluded, obstacle-dense environment using the proposed occlusion-aware CMPC. The occlusion-aware CMPC generates three trajectories (orange, red, and blue) with different considerations of risk regions. All trajectories share an initial consensus segment (green) to enable smooth transitions between trajectories and ensure motion consistency.
@@
To ensure safe navigation in obstacle-dense environments, model predictive control (MPC) facilitates obstacle avoidance through receding-horizon optimization, which systematically accounts for system dynamics and environmental constraints to achieve proactive collision avoidance [15, 13]. To construct obstacle avoidance constraints, accurate prediction of obstacle states is essential for this strategy. Hence, robots must account for potential states of occluded obstacles to guarantee safety in occluded scenarios [8]. However, precise state estimation of fully occluded obstacles remains infeasible [26]. An alternative way involves predicting the risk regions of these obstacles, which reduces planning complexity and enhances safety margins [24]. For instance, OA-MPC [8] adopts worst-case assumptions to ensure safety. This approach theoretically covers potential states of occluded obstacles. However, the generated trajectory tends to be overly conservative [25]. Moreover, a large deviation from the reference path compromises motion consistency [33].
@@
To enhance motion consistency, recent studies have proposed scenario-adaptive planning strategies [6, 35]. In these approaches, multiple homotopic locally optimal trajectories are generated for autonomous vehicles. Although consistent parameters are used within the decision-making module to select the optimal trajectory and maintain motion consistency, each trajectory is optimized separately. Essentially, the lack of coordination can lead to velocity fluctuations, particularly in occluded, obstacle-dense settings. Alternatively, branch MPC (BMPC) [16, 32] introduces a shared common consensus segment across branches, ensuring smooth transitions between trajectory branches and maintaining motion consistency. Similarly, partially observable Markov decision processes (POMDP) [14] ensures environmental consistency through belief state updates [12]. However, these approaches encounter scalability challenges, as branch proliferation significantly increases the number of optimization variables, leading to intractable computation. To address this challenge, recent works have employed the alternating direction method of multipliers (ADMM), which decomposes the original problem into low-dimensional sub-problems to enable real-time performance without compromising safety [34, 21, 10]. Nonetheless, despite the strong computational performance of ADMM, a more definite strategy is still lacking on how to extend its usage to address the challenges posed by occluded, obstacle-dense environments.
@@
In this study, we introduce a novel occlusion-aware Consistent Model Predictive Control (CMPC) strategy for mobile robots navigating in occluded, obstacle-dense environments. The CMPC optimizes several locally optimal trajectories in parallel to account for potential occlusion risks, as shown in Fig. 1. Subsequently, an ADMM-based optimization strategy is used to decompose the optimization problem into parallel sub-problems to facilitate computational efficiency. The main contributions of this paper are summarized as follows:
@@ II-A Reachability Analysis
Reachability analysis is an efficient approach in occlusion-aware planning, as demonstrated in extensive studies [27, 24, 20, 22]. It creates a reachable set of agents in the occluded regions to construct constraints for planning. In [4], the impact of environmental visibility on mobile robot navigation is considered to ensure safety. An MPC strategy is proposed to apply risk fields from reachable sets for automated vehicles [29]. However, reachability analysis typically focuses on worst-case scenarios, leading to overly conservative behaviors. Additionally, predicting obstacle states based on rules of on-road scenarios [27, 24] is not applicable to occluded, obstacle-dense environments encountered by mobile robots.
@@ II-B POMDP
POMDP has been widely adopted in occlusion-aware planning for its ability to handle partially observable scenarios [14, 18]. It leverages a probability distribution in the prediction of the agents’ future states for effective planning. For instance, works [7, 31] leverage the POMDP framework to address navigation and decision-making problems in dynamic environments with partial occlusions. [12] further incorporates a belief state updating module to predict the scenarios more precisely for effective planning. While these approaches can tackle partially observed obstacles in structured road environments, their high computational complexity poses challenges for real-time planning. Additionally, obtaining reliable probability distributions is difficult in unstructured environments with dense occlusion, potentially leading to planning failures.
@@ II-C Contingency Planning
Contingency planning enhances safety and motion consistency by providing multiple contingency paths to address uncertainties in the future states of dynamic obstacles [11, 3, 16, 30]. [23] presents a learning-based method for predicting the states of partially observed agents and motion planning on road. However, existing approaches typically require substantial computational resources to determine policies based on assumptions about future states. Additionally, it typically makes rule-based assumptions about the future states in on-road scenarios, or one-to-one interaction between agents [33]. These future state assumptions may not be reliable in general scenarios.
In this study, we follow the concept of reachability analysis to model the risk regions where occluded obstacles may appear in the future. We predict multiple configurations of risk regions while avoiding explicit hypotheses about obstacle states. To address computational complexity, we propose an occlusion-aware CMPC strategy combined with an ADMM-based optimization process. This strategy ensures safety, efficiency, and adaptability in occluded, obstacle-dense environments.
@@ III Problem Statement
where $k$ denotes the time step; and $\Delta t$ denotes the discrete time step. $f(s(k),u(k),\Delta t)$ is defined as follows:
@@
When an obstacle’s center is not visible in the robot’s FoV, it is treated as an occluded obstacle.
@@
The objective of this work is to develop a motion planning strategy that generates a set of trajectories, enabling safe and consistent navigation in occluded, obstacle-dense environments. The proposed strategy must satisfy the following criteria when navigating in occluded, obstacle-dense environments:
Computational efficiency: Ensure real-time parallel trajectory generation.
Safety guarantee: Guarantee collision-free navigation.
@@ IV-A Risk Regions Modeling
To address the uncertainty of the states of occluded obstacles, we dynamically model risk regions to represent potential locations of occluded obstacles.
@@ IV-A1 Definition of Occluded Regions
Occluded regions are defined as areas blocked by obstacles within the robot’s FoV. For each occluded region, in the robot’s body frame with its center as the origin, the occluded region is defined as the set of points that satisfy the following conditions:
@@
Let $m_{1}$ and $m_{2}$ denote the slopes of the two tangent lines that form the boundaries of the occluded region, $x_{\text{rel}}$ and $y_{\text{rel}}$ denote the position of the obstacle in the robot’s body frame, and $r_{\text{obs}}$ denotes the radius of the obstacle. The expressions for the two tangent lines are defined as follows:
@@
The occluded region $\mathcal{C}$ is represented by the set of points $\bm{x}=\begin{bmatrix}x\\ y\end{bmatrix}$ in the 2D plane that satisfy the following condition:
$\mathcal{C}=\{\bm{x}\in\mathbb{R}^{2}|\bm{A}\bm{x}<0\},$ (6)
where $\bm{A}=\begin{bmatrix}-m_{1}&1\\ m_{2}&-1\\ -1&0\end{bmatrix}$ denotes the coefficient matrix that represents the boundaries of the occluded region.
Figure 2: Modeling of occluded regions and risk regions. Occluded regions are bounded by two tangent lines. Risk regions are located along two tangent lines, each with center $(c_{x},c_{y})$ and radius $r_{\text{risk}}$.
@@ IV-A2 Definition of Risk Regions
Based on the defined occluded regions, we further delineate several risk regions that affect the robot’s motion. Typically, obstacles at the edge of the occluded regions pose the greatest safety threat to the robot. Therefore, these risk regions are defined along the two tangent lines that form the occluded region described in Section IV-A1. Specifically, to provide an effective geometric approximation of the polygonal risk cones along the tangent lines, we model two circular risk regions on each tangent line, as shown in Fig. 2. For the $i$-th circular risk region, the center $(c_{x}^{(i)},c_{y}^{(i)})$ is defined as follows:
$\begin{bmatrix}c_{x}^{(i)}\\ c_{y}^{(i)}\end{bmatrix}$= $\left(id_{\text{risk}}+\sqrt{d_{\text{obs}}^{2}-r_{\text{obs}}^{2}}\right)\cdot\bm{R}\,\frac{1}{\sqrt{1+m_{j}^{2}}}\begin{bmatrix}1\\ m_{j}\end{bmatrix}$ (7)
$+\begin{bmatrix}p_{x}\\ p_{y}\end{bmatrix},\quad j\in\mathcal{I}_{1}^{2},\ i\in\mathcal{I}_{0}^{N_{\text{risk}}-1},$
where $m_{j}$ denotes the slope of the $j$-th tangent line of the occluded region; $d_{\text{risk}}$ denotes the distance between consecutive risk regions; $d_{\text{obs}}$ denotes the distance between the robot and the obstacle; $N_{\text{risk}}$ denotes the total number of risk regions in each configuration; and $\mathcal{I}_{a}^{b}$ denotes a sequence of integers from $a$ to $b$. Since the tangent lines are defined in the robot’s body frame, we transform them to the global frame via a rotation matrix $\bm{R}$.
@@
$r_{\text{risk}}^{(i)}=\frac{\left\|\,(c_{x}^{(i)},c_{y}^{(i)})-(p_{x},p_{y})\,\right\|_{2}}{v+\sigma}\cdot v_{\text{obs,max}}+r_{\text{obs}},$ (8)
where $\sigma\in\mathbb{R}^{+}$ is a small regularization constant (e.g., $\sigma=10^{-4}$) that ensures numerical stability; $v$ denotes the velocity of the robot; and $v_{\text{obs,max}}$ denotes the assumed maximum velocity of the occluded obstacle.
In this study, the conservatism in trajectory planning can be adjusted via different assumptions about the maximum velocity $v_{\text{obs,max}}$ to generate a sequence of $N_{z}$ risk region configurations.
@@
By utilizing the maximum possible obstacle velocity $v_{\text{obs,max}}$ in (8), our circular risk region model provides a conservative over-approximation that encompasses potential positions an occluded obstacle may reach during the planning horizon. For non-circular obstacles, we use their minimum bounding circles to define the risk regions, ensuring the generality of our formulation.
@@ IV-B Occlusion-Aware CMPC
To complete the navigation task and enhance motion consistency, we define a cost function $\mathcal{J}$ over a planning horizon with $N$ time steps to address these requirements:
@@
where $C_{\text{acc}}(\cdot)$ regulates the robot’s acceleration; $C_{\text{vel}}(\cdot)$ represents the velocity tracking cost, measured by the velocity error between the robot and reference velocity; and $C_{\text{guide}}(\cdot)$ represents the deviation from the guidance points. $w_{\text{acc}},\ w_{\text{vel}},\ w_{\text{guide}}\in\mathbb{R}^{+}$ are the corresponding weights for the cost components.
@@
A set of guidance points is generated using a Visual-PRM approach [5]. It produces optimal guidance paths for the receding horizon based on the global task, considering all visible obstacles. Specifically, guidance points provide the endpoints of each horizon, so the cost $C_{\text{guide}}(\cdot)$ is only calculated at the terminal time step k=N1k=N-1.
@@
$s(k)=\tilde{s}(k),\quad\forall k\in\mathcal{I}_{0}^{N_{c}-1},$ (10e)
@@
$H_{\text{kin}}(s)=s(k+1)-f(s(k),u(k),\Delta t),$ (11a)
$\forall k\in\mathcal{I}_{0}^{N-1},$ (11b)
@@
$\forall\{(x_{\text{obs}},y_{\text{obs}}),r_{\text{obs}}\}\in\{\mathcal{B}^{(j)}\}_{j=0}^{N_{\text{obs}}-1},\ \forall s\in\mathcal{S},$ (11c)
@@
$\forall\{(c_{x}^{(i)},c_{y}^{(i)}),r_{\text{risk}}^{(i)}\}\in\{\mathcal{R}^{(i)}_{z}\}_{i=0}^{N_{\text{risk}}-1},\ \forall s\in\mathcal{S}.$ (11d)
Here, $H_{\text{kin}}(\cdot)$ in (11b) enforces kinematic constraints derived from the robot’s kinematic model defined in (3); $G_{\text{obs}}(\cdot)$ in (11c) ensures a minimum safety distance $r_{\text{obs}}$ between the robot and obstacle centers to avoid collision; $\mathcal{B}^{(j)}$ represents the state of the $j$-th obstacle, with the position $(x_{\text{obs}},y_{\text{obs}})$ and the radius of the obstacle $r_{\text{obs}}$; $N_{\text{obs}}$ denotes the number of obstacles considered; $G_{\text{risk}}(\cdot)$ in (11d) maintains a risk boundary $r_{\text{risk}}^{(i)}$ around risk regions; $\mathcal{R}^{(i)}_{z}$ represents the states of the $i$-th risk region in the $z$-th configuration, with center and radius derived from (7) and (8); $N_{z}$ denotes the total number of risk region configurations; (10e) represents the consensus constraint, with consensus variable $\tilde{s}$; and $N_{c}$ denotes the number of time steps in the consensus segment.